The current asset section of Guardian Consultant’s balance sheet consists of cash, accounts receivable, and prepaid expenses. The 2021 balance sheet reported the following: cash, $1,460,000; prepaid expenses, $520,000; long-term assets, $2,500,000; and shareholders’ equity, $2,600,000. The current ratio at the end of the year was 2.5 and the debt to equity ratio was 1.5. Required: Determine the following 2021 amounts and ratios: (Round your "The acid-test ratio" answer to 1 decimal place.)

current liabilities?

Long term liabilities?

Accounts Receivable?

The acid-test ratio?
Answer #1

Answer:

(i)
Debt to equity ratio = Total liabilities/Equity
1.5 = Total liabilities/2,600,000
Total liabilities = $3,900,000
(ii)
Total assets = Total liabilities + Equity
= 3,900,000 + 2,600,000
= $6,500,000
(iii)
Total assets = Current assets + Long term assets
6,500,000 = Current assets + 2,500,000
Current assets = $4,000,000
(iv)
Current assets = Cash + Accounts receivable + Prepaid expenses
4,000,000 = 1,460,000 + Accounts receivable + 520,000
Accounts receivable = $2,020,000
(v)
Current ratio = Current assets/Current liabilities
2.5 = 4,000,000/Current liabilities
Current liabilities = $1,600,000
(vi)
Total liabilities = Current liabilities + Long term liabilities
3,900,000 = 1,600,000 + Long term liabilities
Long term liabilities = $2,300,000
(vii)
Acid test ratio = (Cash + Accounts receivable)/Current liabilities
= (1,460,000 + 2,020,000)/1,600,000
=3,480,000/1,600,000
= 2.175
1 Current liabilities $1,600,000
2 Long term liabilities $2,300,000
3 Accounts receivable $2,020,000
4 Acid test ratio 2.175